2. Suppose a population P(t) satisfies the logistic differential equation dP dt = 0.1P 1 − P 2000 P(0) = 100 Find the following: a) P(20) b) When will the population reach 1200?2. Suppose a population P(t) satisfies the logistic differential equation 2P = 0.1P (1–2000) = 0.1P | P(0) = 100 2000 Find th
Answer #1

del = 0.1811-2000) a = 0-1P ( 200000) Sepasating the vanables - a P = 0ol at f (2000-P) 200097 Resolving into pashal frachons

= lot + 2000-P In poop = lot + in 1 @ P(20) at t= 20 in P a = 10 (20) 2000-p (20) + Incan) in P = 2000 & In 19 2000-P e 200 i